Xcube-x+1ka whole square= 2001 then find value of x which is present in this math question

Grade:10

1 Answers

Arun
6 years ago
Dear Saket
 
x^3 – (x + 1)^2 = 2001
 
x^3 – x^2 – 2x = 2002
x (x^2 – x – 2) = 2002
 
x * (x – 2) (x + 1 ) = 14 * 11 * 13
 
therefore x = 13